Session: Mar 2010 (session 1)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

JIVE (Jun Yang at Mon Feb 15 12:16:11 2010) Success :
Please upload your rxg file once the calibration experiment is finished.

Onsala (Michael Lindqvist at Thu Mar 04 16:11:22 2010) Success:
No known problems. Apart from a lot of RFI :-(.

Medicina (G.Maccaferri at Thu Mar 04 16:36:44 2010) Success:

Shanghai (Bo Xia at Thu Mar 04 16:47:23 2010) Success:

Noto (P. Cassaro at Thu Mar 04 18:17:49 2010) Failure: WIND
Not completely failed, but there is no choice for "50/50". The fringes have been found, but the experiment has been stopped at 13:45 UT due to very strong wind.

Torun (Kaz Borkowski at Fri Mar 05 06:44:07 2010) Minor failures: RECO
A problem with Mk5 prevented us from timely start. Good recordings begun 16 minutes later. Otherwise no known peroblems.

Westerbork (Geert Kuper at Fri Mar 05 08:57:27 2010) Success:
12 Telescopes in the array (no RT1 and RT5) No known problems

Effelsberg (dg at Mon Mar 08 08:53:51 2010) Minor failures: RECO
mk5 grabbed by other terminal: start late at 12:39

Jodrell_Bank (alastair gunn at Mon Mar 29 10:02:54 2010) Minor failures: ANT
Frequent telescope encoder faults resulting in brief periods off source. No other known problems.

Cambridge (alastair gunn at Mon Mar 29 10:05:42 2010) Minor failures:RECE
Faulty final LO synthesiser replaced about 14:15 UT affecting Cm but not Kn. Kn on BBCs 1 and 2. No other known problems.